# Say My Name: a Model’s Bias Discovery Framework

### v1

@@ Appendix E Outputs of SaMyNa
In this section, we provide the detailed output of our bias naming process for the experiments described in the main paper. We report the obtained keywords alongside their associated cosine similarity, in decreasing order of value, for Waterbirds (Tab. 6), CelebA (Tab. 7), BAR (Tab. 8), ImageNet-A (Tab. 9 and 10), on a completely balanced version of Waterbirds (Tab. 11) and for the ablation study on $k$ (Tab. 12).
@@
tree 0.37119 sea 0.55190
forest 0.36188 ocean 0.53744
trees 0.35160 beach 0.43492
forested 0.35120 waters 0.41526
foliage 0.34351 shoreline 0.39085
branch 0.31807 shore 0.37826
stalks 0.31405 coastal 0.37666
vegetation 0.26898 water 0.31551
leaves 0.23139 boat 0.29067
midflight 0.28737
waves 0.27717
seagull 0.26215
lagoon 0.24614
flight 0.22423
lake 0.20760
Table 6: Similarity scores for the Waterbirds dataset.
@@
man 0.40964 woman 0.48667
male 0.38960 makeup 0.25876
lipstick 0.23692
eyeshadow 0.20672
@@
Climbing Diving Fishing Racing Throwing Vaulting
cliff 0.52706 scuba 0.57237 boat 0.45474 cars 0.37512 pitch 0.55673 midair 0.41947
rock 0.44902 underwater 0.54212 river 0.45005 car 0.35930 baseball 0.52524 jump 0.40536
rocks 0.40822 submerged 0.38070 sea 0.42662 track 0.32670 pitcher 0.50494 pole 0.33902
steep 0.35067 coral 0.28247 ocean 0.38038 stadium 0.26144 batter 0.48165 high 0.31889
backpack 0.32901 depths 0.27969 lake 0.35662 speeds 0.24373 player 0.47083 suspended 0.31659
rocky 0.32047 ocean 0.25977 water 0.34019 batter 0.23747 mound 0.42858 athleticism 0.31222
rugged 0.25324 sea 0.25319 marine 0.29564 competitive 0.23413 glove 0.41766 bar 0.27459
jagged 0.24785 wetsuit 0.24872 waves 0.27669 speed 0.22976 athlete 0.40207 outstretched 0.25279
adventurous 0.21917 swimming 0.24449 coral 0.24126 competition 0.22619 playing 0.39830 agility 0.24972
trail 0.21789 marine 0.23934 underwater 0.23954 asphalt 0.22051 sports 0.39558 upward 0.24567
strength 0.21506 water 0.22909 wetsuit 0.21871 baseball 0.21841 sport 0.38070 athletic 0.23762
exploring 0.21200 rocks 0.21742 pitch 0.21645 field 0.36654 athlete 0.22750
nature 0.21199 divers 0.21735 team 0.20312 elbow 0.36480 casting 0.20631
ascending 0.21090 rocky 0.20966 uniform 0.20134 ball 0.35208 prowess 0.20590
adventure 0.20970 scuba 0.20206 pitcher 0.20080 athletic 0.35054 highstakes 0.20370
expedition 0.20443 game 0.34915 arched 0.20291
team 0.31384 sky 0.20140
athleticism 0.30591
arm 0.28327
stadium 0.27959
jersey 0.26615
main 0.25232
action 0.25062
arms 0.24658
activity 0.24016
striking 0.23857
outstretched 0.23707
position 0.23591
uniform 0.23211
serving 0.22028
actions 0.21900
focused 0.21649
skill 0.21620
foreground 0.21460
center 0.21259
stands 0.21227
emphasizes 0.20999
physical 0.20860
catch 0.20644
emphasizing 0.20625
overall 0.20215
agility 0.20136
Table 8: Similarity scores for the BAR dataset.
@@
plumage 0.21741 seagull 0.25743
feathers 0.20849 coastal 0.24091
sea 0.22931
@@
tree: $0.27\pm 0.01$ cranes: $0.34\pm 0.01$ foliage: $0.37\pm 0.02$ ocean: $0.46\pm 0.01$ tree: $0.37\pm 0.01$ sea: $0.54\pm 0.01$ tree: $0.36\pm 0.01$ sea: $0.58\pm 0.02$ foliage: $0.37\pm 0.00$ sea: $0.56\pm 0.01$
branches: $0.21\pm 0.00$ lake: $0.31\pm 0.02$ shrubs: $0.35\pm 0.02$ beach: $0.37\pm 0.01$ forest: $0.35\pm 0.02$ ocean: $0.51\pm 0.00$ forest: $0.36\pm 0.01$ ocean: $0.56\pm 0.02$ forest: $0.37\pm 0.01$ ocean: $0.55\pm 0.01$
brown: $0.19\pm 0.13$ lagoon: $0.31\pm 0.01$ forest: $0.34\pm 0.03$ water: $0.34\pm 0.01$ trees: $0.33\pm 0.03$ beach: $0.44\pm 0.02$ foliage: $0.36\pm 0.01$ beach: $0.49\pm 0.02$ trees: $0.33\pm 0.01$ beach: $0.50\pm 0.01$
trees: $0.18\pm 0.13$ water: $0.27\pm 0.01$ stalks: $0.33\pm 0.02$ coastal: $0.32\pm 0.02$ branch: $0.33\pm 0.00$ coastal: $0.38\pm 0.01$ trees: $0.33\pm 0.01$ coastal: $0.43\pm 0.02$ branch: $0.31\pm 0.00$ shore: $0.44\pm 0.00$
colors: $0.17\pm 0.12$ sea: $0.26\pm 0.18$ tree: $0.33\pm 0.03$ lake: $0.30\pm 0.02$ vegetation: $0.27\pm 0.02$ water: $0.34\pm 0.02$ branch: $0.31\pm 0.01$ water: $0.37\pm 0.02$ vegetation: $0.29\pm 0.00$ water: $0.36\pm 0.01$
leaves: $0.16\pm 0.12$ pier: $0.25\pm 0.18$ trees: $0.32\pm 0.03$ shore: $0.26\pm 0.18$ leaves: $0.24\pm 0.03$ waves: $0.28\pm 0.00$ vegetation: $0.27\pm 0.01$ waves: $0.32\pm 0.01$ leaves: $0.25\pm 0.01$ waves: $0.33\pm 0.01$
black: $0.13\pm 0.10$ shoreline: $0.25\pm 0.17$ vegetation: $0.28\pm 0.02$ waves: $0.24\pm 0.00$ shrubs: $0.23\pm 0.16$ shoreline: $0.28\pm 0.20$ plants: $0.21\pm 0.15$ lake: $0.29\pm 0.02$ shrubs: $0.24\pm 0.17$ lake: $0.27\pm 0.01$
foliage: $0.09\pm 0.13$ crane: $0.21\pm 0.00$ leaves: $0.28\pm 0.02$ pond: $0.23\pm 0.16$ foliage: $0.22\pm 0.15$ midflight: $0.28\pm 0.00$ forested: $0.11\pm 0.15$ seagull: $0.27\pm 0.01$ plants: $0.22\pm 0.15$ midflight: $0.27\pm 0.01$
forests: $0.09\pm 0.12$ ship: $0.15\pm 0.21$ branch: $0.27\pm 0.01$ lagoon: $0.19\pm 0.13$ stalks: $0.21\pm 0.15$ shore: $0.27\pm 0.19$ branches: $0.10\pm 0.14$ midflight: $0.26\pm 0.01$ forested: $0.11\pm 0.15$ seagull: $0.15\pm 0.11$
seagull: $0.15\pm 0.10$ plants: $0.22\pm 0.15$ sea: $0.16\pm 0.22$ forested: $0.12\pm 0.17$ lake: $0.27\pm 0.03$ branches: $0.10\pm 0.14$ shore: $0.16\pm 0.22$ branches: $0.10\pm 0.14$ shoreline: $0.15\pm 0.21$
ocean: $0.13\pm 0.18$ greenishbrown: $0.16\pm 0.12$ seagull: $0.15\pm 0.11$ greenery: $0.08\pm 0.11$ boat: $0.22\pm 0.15$ pond: $0.11\pm 0.15$
vessel: $0.12\pm 0.18$ greenery: $0.16\pm 0.11$ shoreline: $0.14\pm 0.19$ greenishbrown: $0.08\pm 0.11$ pond: $0.21\pm 0.15$

### v2

@@ Appendix G Outputs of SaMyNa
In this section, we provide the detailed output of our bias naming process for the experiments described in the main paper. We report the obtained keywords alongside their associated cosine similarity, in decreasing order of value, for Waterbirds (Tab. 21), CelebA (Tab. 22), BAR (Tab. 25), ImageNet-A (Tab. 23 and 24), on a completely balanced version of Waterbirds (Tab. 26) and for the ablation study on $k$ (Tab. 27).
@@
forest 0.335 ocean 0.540
foliage 0.334 beach 0.434
tree 0.327 shore 0.386
stalks 0.314 water 0.386
branch 0.312 waves 0.335
trees 0.299 boat 0.294
branches 0.296 river 0.268
vegetation 0.273 midflight 0.244
lake 0.244
flying 0.218
Table 21: Cosine similarity scores for keywords associated with the Landbirds and Waterbirds classes.
@@
man 0.45820 woman 0.39947
male 0.39175 blonde 0.34301
mascara 0.24959
makeup 0.24877
lipstick 0.22272
eyeliner 0.20065
@@
Climbing Diving Fishing Racing Throwing Vaulting
climber 0.55677 scuba 0.58300 boat 0.43818 racing 0.42294 pitch 0.54658 vaulter 0.57871
cliff 0.54837 underwater 0.56077 fishing 0.42546 cars 0.42024 baseball 0.48251 vaulting 0.50903
climbing 0.51882 diver 0.49010 river 0.42202 car 0.38508 pitcher 0.48111 vault 0.47589
climb 0.50305 diving 0.34910 ocean 0.39295 track 0.31835 batter 0.46329 jump 0.42575
rock 0.41596 dive 0.33087 lake 0.32940 stadium 0.28773 mound 0.44152 midair 0.37533
steep 0.36449 ocean 0.28606 water 0.31149 race 0.27802 player 0.42868 pole 0.36865
ascent 0.32149 depths 0.27261 marine 0.29890 speeds 0.26571 throw 0.37765 athleticism 0.32692
rocky 0.30365 swimming 0.26935 waves 0.25389 speed 0.26221 athlete 0.37310 suspended 0.28935
ruggedness 0.28319 wetsuit 0.26063 underwater 0.23570 asphalt 0.26199 sport 0.36638 diving 0.27763
jagged 0.26699 marine 0.25820 wetsuit 0.21951 football 0.25381 playing 0.36413 bar 0.26088
ascending 0.20509 water 0.22338 scuba 0.20644 competition 0.23812 glove 0.36378 high 0.24619
pitch 0.23165 elbow 0.34375 fields 0.23763
baseball 0.22954 field 0.33142 arched 0.23102
batter 0.22345 football 0.31316 raised 0.22868
player 0.21114 game 0.31113 athlete 0.22734
team 0.21087 ball 0.30598 dive 0.22547
pitcher 0.20611 team 0.29213 flipper 0.21940
arm 0.28914 skill 0.21579
athleticism 0.27865 peak 0.20534
fields 0.25888
skill 0.23882
striking 0.22815
target 0.22690
focused 0.22104
main 0.21313
position 0.21137
action 0.21031
stadium 0.21019
arms 0.20977
positions 0.20790
pole 0.20539
activity 0.20472
casting 0.20445
stands 0.20009
Table 25: Similarity scores for the BAR dataset.
@@
- - sea 0.31901
midflight 0.23266
seagull 0.22417
@@
midair: $0.18\pm 0.12$ water: $0.24\pm 0.00$ tree: $0.33\pm 0.02$ water: $0.38\pm 0.02$ forest: $0.36\pm 0.02$ ocean: $0.54\pm 0.02$ tree: $0.36\pm 0.01$ ocean: $0.58\pm 0.01$ tree: $0.38\pm 0.02$ ocean: $0.56\pm 0.01$
soars: $0.15\pm 0.11$ seagull: $0.17\pm 0.12$ forest: $0.33\pm 0.03$ beach: $0.37\pm 0.03$ foliage: $0.35\pm 0.01$ beach: $0.44\pm 0.04$ forest: $0.34\pm 0.01$ beach: $0.51\pm 0.02$ forest: $0.37\pm 0.01$ beach: $0.52\pm 0.01$
tree: $0.07\pm 0.10$ fish: $0.17\pm 0.12$ foliage: $0.31\pm 0.02$ seagull: $0.31\pm 0.03$ branch: $0.34\pm 0.02$ water: $0.37\pm 0.02$ foliage: $0.34\pm 0.01$ water: $0.39\pm 0.02$ foliage: $0.35\pm 0.01$ water: $0.37\pm 0.02$
boathouse: $0.17\pm 0.12$ branches: $0.31\pm 0.02$ lake: $0.31\pm 0.02$ trees: $0.33\pm 0.02$ sea: $0.36\pm 0.26$ trees: $0.33\pm 0.01$ lake: $0.30\pm 0.01$ branch: $0.35\pm 0.02$ waves: $0.34\pm 0.02$
seabirds: $0.15\pm 0.10$ trees: $0.30\pm 0.01$ waves: $0.27\pm 0.01$ branches: $0.32\pm 0.02$ waves: $0.33\pm 0.01$ branch: $0.31\pm 0.02$ midflight: $0.23\pm 0.02$ trees: $0.34\pm 0.01$ lake: $0.27\pm 0.02$
ships: $0.14\pm 0.19$ branch: $0.30\pm 0.02$ midflight: $0.27\pm 0.01$ stalks: $0.32\pm 0.01$ boat: $0.31\pm 0.01$ branches: $0.30\pm 0.01$ waves: $0.23\pm 0.16$ leaves: $0.24\pm 0.01$ midflight: $0.26\pm 0.00$
vessel: $0.14\pm 0.19$ stalks: $0.28\pm 0.02$ flying: $0.17\pm 0.12$ tree: $0.24\pm 0.17$ shoreline: $0.27\pm 0.19$ leaves: $0.24\pm 0.01$ sea: $0.20\pm 0.28$ branches: $0.21\pm 0.15$ flying: $0.22\pm 0.01$
ship: $0.13\pm 0.19$ bamboo: $0.26\pm 0.02$ ocean: $0.17\pm 0.24$ vegetation: $0.19\pm 0.13$ shore: $0.26\pm 0.18$ seagull: $0.14\pm 0.10$ sandy: $0.15\pm 0.11$
pier: $0.10\pm 0.15$ vegetation: $0.15\pm 0.11$ sea: $0.16\pm 0.23$ leaves: $0.16\pm 0.11$ lake: $0.26\pm 0.01$ horizon: $0.14\pm 0.10$ horizon: $0.07\pm 0.10$
lake: $0.09\pm 0.13$ leaves: $0.15\pm 0.11$ watermark: $0.16\pm 0.11$ midflight: $0.24\pm 0.02$ rocky: $0.07\pm 0.09$
sea: $0.09\pm 0.13$ deciduous: $0.09\pm 0.12$ shore: $0.12\pm 0.18$ flying: $0.14\pm 0.10$
crane: $0.07\pm 0.10$ stalk: $0.08\pm 0.11$ river: $0.09\pm 0.13$ coastal: $0.11\pm 0.16$
dock: $0.07\pm 0.10$ seabird: $0.08\pm 0.12$ river: $0.09\pm 0.13$
horizon: $0.07\pm 0.10$ foam: $0.09\pm 0.13$